OpenStudy (anonymous):

"Write each expression as the product of two binomials" 2x^3-3x^2-4x+6.

OpenStudy (raden):

you want factor out it , right ? 2x^3-3x^2-4x+6. it can be written as : (2x^3 - 3x^2) + (-4x + 6) look at in the first paratheses, the common factor is x^2 and the 2nd parantheses has the common factor is -2. therefore, it can be (2x^3 - 3x^2) + (-4x + 6) = x^2(2x - 3) - 2(2x - 3) and finally, by using the distributive property we have the product of 2 binomials : x^2(2x - 3) - 2(2x - 3) = (2x - 3)(x^2 - 2)

OpenStudy (anonymous):

Hmm. Can you show how to use the distributive property for that 2nd part?

OpenStudy (anonymous):

And yeah, I want to factor it out.

OpenStudy (anonymous):

Oh wait, never-mind, I get it now. Thanks.

OpenStudy (raden):

wlcm :)